# Exact Random Graph Matching with Multiple Graphs

### v1

@@ 4.3 Piecing it all together: Proof of Theorem 3
$\stackrel{{\scriptstyle\text{(c)}}}{{\leq}}\!\sum_{\ell=1}^{\lfloor m/2\rfloor}m^{\ell}\left[o\left(\frac{1}{n}\right)+o\left(\frac{1}{n}\right)\right]=o\left(\frac{1}{n}\right).$
Here, (a) uses Theorem 9, and (b) uses the fact that for any $\ell\geq 2$, the random variable $\widetilde{c}_{v}(U_{\ell})$ stochastically dominates $\widetilde{c}_{v}(U_{1})$ (Theorem 10). Finally, (c) uses Theorem 11 and the fact that $Cs(1-(1-s)^{m-1})>1$. Therefore, a union bound over all the nodes yields
$\mathbb{P}\left(E_{2}\right)\leq o(1)+\sum_{v\in V}q_{v}\leq o(1)+n\times o(1/n)=o(1).$
Finally, by optimality of the MLE, it follows that
$\mathbb{P}\left(E_{1}\right)\leq\mathbb{P}\left(E_{2}\right)=o(1),$
whenever $Cs(1-(1-s)^{m-1})>1$. This concludes the proof. ∎
5 Discussion and Future Work
In this work, we introduced and analyzed matching through transitive closure - an approach that combines information from multiple graphs to recover the underlying correspondence between them. Despite its simplicity, it turns out that matching through transitive closure is an optimal way to combine information in the setting where the graphs are pairwise matched using the $k$-core estimator. A limitation of our algorithms is the runtime: Algorithm 2 does not run in polynomial time because it uses the $k$-core estimator for pairwise matching, which involves searching over the space of permutations. Even so, it is useful to establish the fundamental limits of exact recovery, and serve as a benchmark to compare the performance of any other algorithm.
The transitive closure subroutine (Step 2) itself is efficient because it runs in polynomial time $O(mn)$. Therefore, a natural next step is to modify Step 1 in our algorithm so that the pairwise matchings are done by an efficient algorithm. However, it is not clear if transitive closure is optimal for combining information from the pairwise matchings in this setting. For example, there is a possibility that the pairwise matchings resulting from the efficient algorithm are heavily correlated, and transitive closure is unable to boost them. In Figure 3, we show experimentally that this is not the case for two algorithms of interest: GRAMPA [FMWX22] and Degree Profiles [DMWX21].
@@ 5 Discussion and Future Work
    GRAMPA is a spectral algorithm that uses the entire spectrum of the adjacency matrices to match the two graphs. The code is available in [FMWX20].
@@
    Degree Profiles associates with each node a signature derived from the degrees of its neighbors, and matches nodes by signature proximity. The code is available in [DMWX20].
Evidently, both algorithms benefit substantially from using transitive closure to boost the number of matched nodes. This suggests that transitive closure can be a practical algorithm to boost matchings between networks by using other networks as side-information. Unfortunately, both GRAMPA and Degree Profiles require the graphs to be close to isomorphic in order to perform well, and so they do not perform well when the model parameters are close to the information theoretic threshold for exact recovery. Subsequently, they cannot be used to answer the question in 1.
Our work presents several directions for future research.
Figure 3: Matching through transitive closure
@@
  • Polynomial-time algorithms. Using a polynomial-time estimator in place of the $k$-core estimator in Step 1 of Algorithm 2 yields a polynomial-time algorithm to match $m$ graphs. It is critical that the estimator in question is able to identify for itself the nodes that it has matched correctly - this precision is present in the $k$-core estimator and enables the transitive closure subroutine to work correctly. Can the performance guarantees of the $k$-core estimator be realized through polynomial time algorithms that meet this constraint?
  • Beyond Erdős-Rényi graphs. The study of matching two ER graphs provided tools and techniques that extended to the analysis of more realistic models. For instance, the $k$-core estimator itself played a crucial role in establishing limits to matching two correlated stochastic block models [GRS22] and two inhomogeneous random graphs [RS23]. Can the techniques developed in the present work be used to identify the information theoretic limits to exact recovery in these models in the general setting of $m$ graphs?
@@
  • Robustness. Finally, how sensitive to perturbation is the transitive closure algorithm? Is it possible to quantify the extent to which an adversary may perturb edges in some of the graphs without losing the performance guarantees of the matching algorithm? Algorithms that perform well on models such as ER graphs and are further generally robust are expected to also work well with real-world networks.
Acknowledgments and Disclosure of Funding
This work was supported by NSF under Grant CCF 19-00636.
Appendix A Maximum Likelihood Estimator
Recall the form of the maximum likelihood estimator as claimed in Theorem 4. See 4
Proof.
Notice that

### v2

@@ 5.2 Piecing it together: Proof of Theorem 1
$\stackrel{{\scriptstyle\text{(b)}}}{{\leq}}\!o(1)+\sum_{\ell=1}^{\lfloor m/2\rfloor}\!m^{\ell}\cdot\mathbb{P}\left(\exists\ v\in[n]:\widetilde{c}_{v}(U_{1})\leq\frac{m^{2}}{4}\big{(}k\!+\!2/\alpha\big{)}\!\right)$
$\stackrel{{\scriptstyle\text{(c)}}}{{\leq}}\!o(1)+\sum_{\ell=1}^{\lfloor m/2\rfloor}m^{\ell}\cdot\mathbb{P}\left(\exists\ v\in[n]:\big{\{}\delta_{G_{1}\wedge G_{2}^{\prime}}(v)\leq r\big{\}}\cap\cdots\cap\big{\{}\delta_{G_{1}\wedge G_{m}^{\prime}}(v)\leq r\big{\}}\right)\stackrel{{\scriptstyle\text{(d)}}}{{\leq}}o(1)$
Here, (a) uses Lemma 24 and a union bound, and (b) uses the fact that for any $\ell\geq 2$, the random variable $\widetilde{c}_{v}(U_{\ell})$ stochastically dominates $\widetilde{c}_{v}(U_{1})$ (Lemma 25). Lastly, (c) uses the definition of $\widetilde{c}_{v}(U_{1})$ and (d) uses Lemma 26. Therefore, transitive closure successfully matches all nodes in $[n]$ with probability $1-o(1)$. It follows that $\mathbb{P}\left(E_{2}\right)=o(1)$. This concludes the proof. $\Box$
5.3 Proofs of Lemmas 24 and 26
5.3.1 Proof of Lemma 24.
For any vertex cut $U$,
$\left\{\widetilde{c}_{v}(U)>\frac{m^{2}}{4}\left(k+\frac{2}{\alpha}\right)\right\}$ $\stackrel{{\scriptstyle\text{(a)}}}{{\subseteq}}\left\{\widetilde{c}_{v}(U)>|U|\left(m-|U|\right)\left(k+\frac{2}{\alpha}\right)\right\}$
=$=\bigg{\{}\sum_{i\in U}\sum_{j\in U^{c}}\delta_{G_{i}^{\prime}\wedge G_{j}^{\prime}}(v)>|U|(m-|U|)\bigg{(}k+\frac{2}{\alpha}\bigg{)}\bigg{\}}$
$\subseteq\bigcup_{i\in U}\bigcup_{j\in U^{c}}\left\{\delta_{G_{i}^{\prime}\wedge G_{j}^{\prime}}(v)>k+\frac{2}{\alpha}\right\},$
where (a) uses the fact that the maximum of a set of numbers is greater than or equal to the average. On the other hand,
$\left\{c_{v}(U)=0\right\}=\bigcap_{i\in U}\bigcap_{j\in U^{c}}\left\{v\notin\mathrm{core}_{k}(G_{i}^{\prime}\wedge G_{j}^{\prime})\right\}.$
$p_{1}\leq\sum_{i\in U}\sum_{j\in U^{c}}\mathbb{P}\left(\left\{v\notin\mathrm{core}_{k}(G_{i}^{\prime}\wedge G_{j}^{\prime})\right\}\cap\left\{\delta_{G_{i}^{\prime}\wedge G_{j}^{\prime}}(v)>k+\frac{2}{\alpha}\right\}\right)=o(1/n),$
[…]
5.3.2 Proof of Lemma 26
If $s=1$, then the graphs $G_{1},\cdots,G_{m}$ are all isomorphic, and
$(\star)=\mathbb{P}\left(\exists v\in V\setminus v^{*}:\delta_{G_{1}}(v)\leq r\right)\leq\sum_{v\in V\setminus v^{*}}\mathbb{P}\left(\delta_{G_{1}}(v)=0\right)\leq\sum_{i=1}^{n}e^{-d_{i}}=o(1),$
and so the result follows. We therefore assume without loss of generality that $s<1$. Observe that
(($\star$)( ⋆ ) $\leq\mathbb{P}\left(\bigcup_{r_{2}=0}^{r}\cdots\bigcup_{r_{m}=0}^{r}\left(\exists\ v\in V\setminus v^{*}:\big{\{}\delta_{G_{1}\wedge G_{2}^{\prime}}(v)=r_{2}\big{\}}\cap\cdots\cap\big{\{}\delta_{G_{1}\wedge G_{m}^{\prime}}(v)=r_{m}\big{\}}\right)\right)$
$\leq\sum_{r_{2}=0}^{r}\cdots\sum_{r_{m}=0}^{r}\ \ \sum_{v\in V\setminus v^{*}}\mathbb{E}_{D_{v}}\left[\mathbb{P}\left(\big{\{}\delta_{G_{1}\wedge G_{2}^{\prime}}(v)=r_{2}\big{\}}\cap\cdots\cap\big{\{}\delta_{G_{1}\wedge G_{m}^{\prime}}(v)=r_{m}\big{\}}\mid\delta_{G_{1}}(v)=D_{v}\right)\right]$
=$\stackrel{{\scriptstyle\text{(a)}}}{{=}}\sum_{r_{2}=0}^{r}\cdots\sum_{r_{m}=0}^{r}\ \ \sum_{v\in V\setminus v^{*}}\mathbb{E}_{D_{v}}\left[\prod_{\ell=2}^{m}\binom{D_{v}}{r_{\ell}}s^{r_{\ell}}(1-s)^{D_{v}-r_{\ell}}\cdot\bm{1}_{\{r_{\ell}\leq D_{v}\}}\right]$
$\stackrel{{\scriptstyle\text{(b)}}}{{\leq}}\sum_{r_{2}=0}^{r}\cdots\sum_{r_{m}=0}^{r}\left[\prod_{\ell=2}^{m}\left(\frac{e}{r_{\ell}}\cdot\frac{s}{1-s}\right)^{r_{\ell}}\right]\cdot\sum_{v\in V\setminus v^{*}}\mathbb{E}_{D_{v}}\left[\prod_{\ell=2}^{m}D_{v}^{r_{\ell}}(1-s)^{D_{v}}\right],$ (25)
where (a) uses that the degrees of node $v$ in the graphs $G_{1}\wedge G_{2}^{\prime},\cdots,G_{1}\wedge G_{m}^{\prime}$ are conditionally independent of the degree of $v$ in $G_{1}$, and that each edge of $G_{1}$ is independently present in $G_{1}\wedge G_{j}^{\prime}$ with probability $s$. Further, (b) uses that $\binom{n}{k}\leq\left(ne/k\right)^{k}$ with the implicit definition $(1/0)^{0}\triangleq 1$. Note that
$\sum_{v\in V\setminus v^{*}}\mathbb{E}_{D_{v}}\left[\prod_{\ell=2}^{m}D_{v}^{r_{\ell}}(1-s)^{D_{v}}\right]$ $=\sum_{v\in V\setminus v^{*}}\left(\sum_{d=0}^{n-1}\mathbb{P}\left(D_{v}=d\right)\prod_{\ell=2}^{m}d^{r_{\ell}}(1-s)^{d}\right)$
$=\sum_{d=0}^{n-1}\left(\sum_{v\in V\setminus v^{*}}\mathbb{P}\left(D_{v}=d\right)\cdot d^{\sum_{\ell=2}^{m}r_{\ell}}\cdot(1-s)^{(m-1)d}\right).$ (26)
Proceed by splitting the outer summation in (26) over $d$ at $d^{*}\triangleq n^{\frac{\alpha}{2(m-1)r}}$. We have
$\sum_{d=0}^{d^{*}}\sum_{v\in V\setminus v^{*}}$ $\mathbb{P}\left(D_{v}=d\right)\cdot d^{\sum_{\ell=2}^{m}r_{\ell}}\cdot(1-s)^{(m-1)d}\leq\sum_{v\in V\setminus v^{*}}(d^{*})^{\sum_{\ell=2}^{m}r_{\ell}}\sum_{d=0}^{d^{*}}\mathbb{P}\left(D_{v}=d\right)\cdot(1-s)^{(m-1)d}$
$\leq(d^{*})^{\sum_{\ell=2}^{m}r_{\ell}}\sum_{v\in V\setminus v^{*}}\mathbb{E}_{D_{v}}\left[(1-s)^{(m-1)D_{v}}\right]$
=$\stackrel{{\scriptstyle\text{(c)}}}{{=}}(d^{*})^{\sum_{\ell=2}^{m}r_{\ell}}\sum_{v\in V\setminus v^{*}}\prod_{u\neq v}\left(1-p_{uv}s+p_{uv}s(1-s)^{m-1}\right)$
$\stackrel{{\scriptstyle\text{(d)}}}{{\leq}}(d^{*})^{\sum_{\ell=2}^{m}r_{\ell}}\sum_{v\in V\setminus v^{*}}\exp\bigg{(}-\sum_{u\neq v}p_{uv}s\left(1-(1-s)^{m-1}\right)\bigg{)}$
$\stackrel{{\scriptstyle\text{(e)}}}{{\leq}}n^{\alpha/2}\cdot\sum_{v\in V\setminus v^{*}}e^{-d_{v}s\left(1-(1-s)^{m-1}\right)}=o(1).$ (27)
Here, (c) uses the moment generating function of $D_{v}$, which is distributed as a sum of independent Bernoulli random variables with means $p_{uv}$ respectively. Further, (d) uses that $1-x\leq e^{-x}$ for any $x$, and finally (e) uses the definition of $d^{*}$ and the condition (24). The other part of the split sum can be bounded as
$\sum_{d=d^{*}}^{n-1}\sum_{v\in V\setminus v^{*}}$ $\mathbb{P}\left(D_{v}\!=\!d\right)\cdot d^{\sum_{\ell=2}^{m}r_{\ell}}\cdot(1-s)^{(m-1)d}\leq\sum_{v\in V\setminus v^{*}}\left[\max_{d\in[d^{*},n]}d^{\sum_{\ell=2}^{m}r_{\ell}}\cdot(1-s)^{(m-1)d}\right]\mathbb{P}\left(D_{v}\geq d^{*}\right)$
$\stackrel{{\scriptstyle\text{(f)}}}{{\leq}}(d^{*})^{\sum_{\ell=2}^{m}r_{\ell}}(1-s)^{(m-1)d^{*}}\cdot\sum_{v\in V\setminus v^{*}}\mathbb{P}\left(D_{v}\geq d^{*}\right)$
$\stackrel{{\scriptstyle\text{(g)}}}{{\leq}}n^{1+\alpha/2}\cdot\left((1-s)^{m-1}\right)^{n^{\frac{\alpha}{2(m-1)r}}}=o(1),$ (28)
whenever $s>0$. Here, (f) uses that the function $d\mapsto d^{\sum r_{i}}(a-s)^{md}$ is decreasing on the interval $[d^{*},n]$ for all sufficiently large $n$. Finally, (g) uses the definition of $d^{*}$ and that $\mathbb{P}\left(D_{v}\geq d^{*}\right)\leq 1$ for all $v$. Plugging (27) and (28) in (26) yields that $\sum_{v\in V\setminus v^{*}}\mathbb{E}_{D_{v}}\left[\prod_{\ell=2}^{m}D_{v}^{r_{\ell}}(1-s)^{D_{v}}\right]=o(1)$. Therefore, continuing from (25),
[…]
$\leq(r+1)(m-1)\cdot\max_{1\leq r^{\prime}\leq r}\left(\frac{es}{r^{\prime}(1-s)}\right)^{r^{\prime}(m-1)}\cdot o(1)=o(1),$
since $r$ and $m$ are both constants independent of $n$. This completes the proof. $\Box$
5.4 Proof of Theorem 5
In this section, we prove Theorem 5, which implies that our sufficient condition (2) is not necessary for matching inhomogeneous graphs in general.
Proof of Theorem 5. The proof is nonconstructive; the idea is to show that recovery is possible if the parameter matrix $P$ is random with a suitable distribution, and then show that a set of deterministic choices for $P$ satisfying the requirements of the proposition has positive probability (actually probability converging to one) under the distribution of $P.$ Let $G_{-1}$ denote an Erdős-Rényi graph with parameters $n$ and $q$, where $q$ is to be determined later. Let $P$ denote the random matrix equal to $p_{\max}$ times the incidence matrix of $G_{-1}.$
Given $P$, let $G\sim\mathrm{RG}(n,P)$, let $G^{\prime}_{1},\cdots G^{\prime}_{m}$ denote the graphs obtained from $G$ by subsampling each edge with probability $s$, and let $G_{k}$ be obtained from $G^{\prime}_{k}$ by randomly uniformly permuting the labels, for $k\in[m]$. We will consider the algorithm that separately matches $G_{-1}$ to $G_{k}$ for each $k$, and then matches $G_{1},\cdots,G_{m}$ to each other via transitive closure through $G_{-1}$. For $k\in[m]$, $G^{\prime}_{k}$ can be viewed as being obtained from $G_{-1}$ by independently subsampling the edges of $G_{-1}$ with probability $p_{\max}s.$ Therefore, if $r_{ij}\triangleq\mathbb{P}\left(\bm{1}_{\{e\in E(G_{-1})\}}=i,\bm{1}_{\{e\in E(G_{k})\}}=j\right)$, then
$\left(\begin{array}[]{cc}r_{00}{}{}&r_{01}\\ r_{10}{}{}&r_{11}\end{array}\right)=\left(\begin{array}[]{cc}1-q&~{}~{}0\\ q(1-p_{\max}s)&~{}~{}qp_{\max}s\end{array}\right)$
It is known (CK, 17; WXY, 22) for correlated Erdős-Rényi graphs such as this that exact recovery is possible if there is some $\varepsilon^{\prime}>0$ such that
$n\left(\sqrt{r_{00}r_{11}}-\sqrt{r_{01}r_{10}}\right)^{2}\geq(1+\varepsilon^{\prime})\log n.$
Assume without loss of generality that $\epsilon<1.$ Let $\epsilon^{\prime}=\epsilon/4$ and select $q$ so that $qs=\frac{1+2\epsilon^{\prime}}{\log n}.$ With that choice we conclude that the algorithm described exactly matches $G_{k}$ to $G_{-1}$ with probability $1-o(1)$ for each $k$, and hence matches $G_{1},\cdots,G_{m}$ to each other with probability $1-o(1).$
Let $\delta_{G_{-1}}(i)$ denote the degree of node $i$ in graph $G_{-1}$. Note that $\delta_{G_{-1}}(i)\sim\text{Bin}(n-1,q)\preceq\text{Bin}(n,q)$ for all nodes $i$. Therefore by the union bound and Chernoff inequality [Theorem 4.4(MU, 17)],
$\mathbb{P}\left(\bigcup_{i=1}^{n}\left\{\delta_{G_{-1}}(i)>(1+\varepsilon^{\prime})nq\right\}\right)\leq n\cdot\mathbb{P}\left(\text{Bin}\left(n,q\right){>}(1+\varepsilon^{\prime})nq\right)\leq ne^{-nq(\varepsilon^{\prime})^{2}/3}=o(1).$ (29)
Let $D_{i}=p_{\max}\cdot\delta_{G_{-1}}(i)$ for $i\in[n],$ so that $D_{i}$ is the random version of $d_{i}.$ Then (3), (29), and the fact $(1+\epsilon^{\prime})(1+2\epsilon^{\prime})<1+\epsilon$ imply that $\mathbb{P}\left(D_{i}\leq d~{}~{}\mbox{for}~{}i\in[n]\right)=1-o(1).$ In summary, the intersection of the following three events has probability $1-o(1):$
@@ 5.4 Proof of Theorem 5
    $\max_{i,j}P_{ij}=p_{\max}=(\log n)^{2}/n$.
@@
    $G_{1},\cdots,G_{m}$ are exactly matched by the transitive closure algorithm via $G_{-1}$.
  3. 3.
    $D_{i}\leq d~{}~{}\mbox{for}~{}i\in[n]$.
[…]
5.5 On Necessary Conditions: Proofs of Theorem 8
In this section, we prove Theorem 8, i.e. our sufficient condition is also necessary for the homogeneous case. Theorem 8 has a simple proof following a genie-aided converse argument: reduce the problem to that of matching two graphs by providing extra information to the estimator.
[…]
6 Simulation Results and Discussion
A result of Rácz and Sridhar (RS, 23) establishes that with high probability, the $k$-core estimator for two inhomogeneous graphs matches only the nodes in the $k$-core of their true intersection graph. This is true even if the matrix $\mathbf{P}$ has constant entries that do not depend on $n$. Although the $k$-core estimator itself is not efficient to implement, its output can be efficiently simulated by computing the $k$-core of the true intersection graph. This is possible only because the simulator has access to the ground truth. Under the assumption that the $k$-core estimator outputs the $k$-core of the true intersection graph, we can simulate the performance of Algorithm 1. Figure 3 shows the mean fraction of matched nodes as a function of $m$ before and after transitive closure, for a variety of graph models. For each value of $m$, the $\binom{m}{2}$ pairs of graphs are pairwise matched by simulating the $k$-core estimator, and the average fraction of matched nodes is recorded. The transitive closure subroutine is then applied to the partial matchings and the improvement in the fraction of matched nodes is recorded. In all figures, the shaded region represents the standard deviation for the fraction of matched nodes. Three settings are studied:
(a) Erdős-Rényi graphs, $(n,p,s,k)=(10^{4},0.003,0.8,13)$
(b) SBM: $(n,p,q,s,k)=(10^{4},0.04,0.01,0.25,14)$
(c) RGG, $(n,p,r,s,k)=(10^{4},0.15,0.2,0.4,14)$
Figure 3: Mean fraction of correctly matched nodes for various graphs obtained by simulating Algorithm 1
  1. 1.
    Erdős-Rényi graphs: Figure 3(a) studies ER graphs on $10^{4}$ nodes. An edge is present in the parent graph with probability $p=0.003$ and subsampled independently to the graphs $G_{1}$ and $G_{2}$ with probability $0.8$.
  2. 2.
    Stochastic Block Model: Figure 3(b) similarly presents the fraction of matched nodes before and after transitive closure for a balanced $5$-community stochastic block model on $10^{4}$ nodes, with intra-community edge probability $p=0.04$ and inter-community edge probability $q=0.01$ respectively. The subsampling probability $s=0.25$.
  3. 3.
    Random Geometric Graph: Figure 3(c) considers random geometric graphs generated as follows: First, $n=10^{4}$ points are sampled independently and uniformly on the unit square $[0,1]\times[0,1]$. The parent graph $G$ is generated by connecting two points within Euclidean distance $r=0.2$ with probability $p=0.15$, and the graphs $G_{1}$ and $G_{2}$ are obtained by subsampling each edge independently with probability $s=0.4$.
All data points are obtained by averaging across $50$ independent runs of the algorithm. As Figure 3 shows, the transitive closure step boosts the fraction of matched nodes across any two graphs.
A limitation of Algorithm 1 is the runtime – it does not run in polynomial time because it uses the $k$-core estimator for pairwise matching, which involves searching over the space of permutations. Even so, it is useful to establish the fundamental limits of exact recovery, and serve as a benchmark to compare the performance of any other algorithm.
The transitive closure subroutine (Step 2) itself is efficient because it runs in polynomial time. Therefore, a natural next step is to modify Step 1 in our algorithm so that the pairwise matchings are done by an efficient algorithm. However, it is not clear if transitive closure is optimal for combining information from the pairwise matchings in this setting. For example, there is a possibility that the pairwise matchings resulting from the efficient algorithm are heavily correlated, and transitive closure is unable to boost them.
7 Conclusion
In this work, we introduced and analyzed matching through transitive closure - a black-box algorithm to combine information from multiple graphs to recover the underlying correspondence between them. Despite its simplicity, it turns out that matching through transitive closure is an optimal way to combine information in the homogeneous setting where the graphs are pairwise matched using the $k$-core estimator. The analysis of this algorithm yielded a sufficient condition to exactly match $m$ inhomogeneous random graphs. Our work presents several directions for future research.
  • Impossibility results for inhomogeneous graphs. The derivation and analysis of the maximum likelihood estimator for exact recovery can help quantify the sharpness of the sufficient condition (2). Can the heterogeneity in the model be exploited to match two graphs even when their true intersection graph is disconnected?
@@ 7 Conclusion
  • Polynomial-time algorithms. Using a polynomial-time estimator in place of the $k$-core estimator in Step 1 of Algorithm 1 yields a polynomial-time algorithm to match $m$ graphs. It is critical that this estimator is precise. Can the performance guarantees of the $k$-core estimator be realized through polynomial time algorithms that meet this constraint?
@@
  • Robustness. Finally, how sensitive to perturbation is the transitive closure algorithm? Is it possible to quantify the extent to which an adversary may perturb edges in some of the graphs without losing the performance guarantees of the matching algorithm? Algorithms that perform well on models such as inhomogeneous graphs and are further generally robust are expected to also work well with real-world networks.
Appendix A Maximum Likelihood Estimator for Exactly Matching Multiple Homogeneous Graphs
We show that in the homogeneous setting, the maximum likelihood estimator is given by the permutation profile that minimizes the number of edges in the corresponding union graph, i.e.
$(\widehat{\pi}^{\text{MLE}}_{12},\cdots,\widehat{\pi}^{\text{MLE}}_{1m})\in\operatorname*{arg\,min}_{\pi_{12},\cdots,\pi_{1m}}|E(G_{1}\vee G_{2}^{\pi_{12}}\vee\cdots\vee G_{m}^{\pi_{1m}})|.$
Lemma 27
Let $G_{1},\cdots,G_{m}$ be correlated Erdős-Rényi graphs obtained from the subsampling model with parameters $p$ and $s$. Further, let $\pi_{12},\cdots,\pi_{1m}$ denote a collection of permutations on $[n]$. Then
$\log\mathbb{P}\left(G_{1},\cdots,G_{m}\mid\pi^{*}_{12}=\pi_{12},\cdots,\pi^{*}_{1m}=\pi_{1m}\right)\propto\mathrm{const}.-|E(G_{1}\vee G_{2}^{\pi_{12}}\vee\cdots\vee G_{m}^{\pi_{1m}})|,$
where $\mathrm{const}.$ depends only on $p$, $s$ and $G_{1},\cdots,G_{m}$.
Proof. Consider the setting where $p_{ij}=p$ for all $i\neq j$. Notice that